OpenStudy (anonymous):

for the natural logs and log base ten, are there any rules like the "stuff rule" that are useful when evaluating or simplifying/converting?

OpenStudy (anonymous):

cause i know that ln e is equal to 1. but what else can it be used for? besides entering into calculator. what's a great use for it? @phi

OpenStudy (phi):

are you asking why we use natural log when we already have log (base 10)? It's helpful in calculus

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wait, you know how you showed me the stuff rule using the base of 10? what if it wasn't the base of 10? cause one problem, i got was 5&(logbase5 8) +3^(logbase3 9)

OpenStudy (anonymous):

5^(logbase5 8) **

OpenStudy (phi):

you undo the log by making it the exponent of the log's base in other words, \[ 5^{\log_5 8} = 8\]

OpenStudy (anonymous):

oh it still works like that? oh nvmd thank you

OpenStudy (phi):

there is one rule which is helpful (though painful to remember) that is the "change of base" formula your calculator only knows base e and base 10 say you want \[ \log_2 8 \] you use this formula \[ \log_2 8 = \frac{\ln 8}{\ln 2} \] or \[ \log_2 8 = \frac{\log 8}{\log 2} \] or using letters \[ \log_a x = \frac{\log_b x}{\log_b a} \]

OpenStudy (anonymous):

oh yeah, my teacher showed me something like that :)